# Explainable Artificial Intelligence techniques for interpretation of food datasets: a review

### v1

@@ 4 Explaining Food Safety
Table 1: Summary of the works introducing applying XAI for food safety surveyed in Section 4, according to their data type and explanation type (labelled as “Expl. type”).
@@
[24] Tabular Mixed
4.1 Visual Explanation
Numerous studies using pictorial data have focused on detecting plant diseases in staple crops like maize, rice, and wheat, showcasing the importance of accurate disease identification in the food supply. [55] developed a transfer learning methodology enhancing MobileNetV2 with CAM to diagnose plant diseases in maize and rice. [66] and [64] used CNN models to detect maize and peanut diseases, applying channel attention and pruning techniques for improved feature extraction. Similarly, [36] applied CNNs with Grad-CAM to distinguish healthy from infected wheat, effectively identifying disease-affected areas. [52] introduced C-DenseNet, a modified CNN model, to grade wheat stripe rust severity, validated using Grad-CAM++ [107]. [56] developed a YOLOv5s-based model with MobileNetV3 and C3Ghost modules to detect fusarium head blight (FHB) in wheat, using Grad-CAM. In addition, addressing the complexities of varying disease images, [37] improved MobileNetV2 with a Location-wise Soft Attention mechanism and CAM, demonstrating its practical utility in identifying crop diseases in diverse conditions.
In contrast, some studies explored alternative crops and agricultural sectors. [72] used high-resolution video data and a CNN-based object detection model to monitor pecan tree health, focusing on xylella disease, validated by Grad-CAM to highlight critical canopy features. [54] introduced T-RNet, a Transformer-Embedded ResNet model, for cassava leaf disease detection, demonstrating focus on relevant areas through Grad-CAM visualizations.
@@ 4.1 Visual Explanation
Similarly, [88] applied LIME and Grad-CAM to a DenseNet-based model developed for classifying tomato leaf diseases. Moreover, [99] proposed the use of a MobileNetV2 model combined with data augmentation and reweighting techniques for accurate classification of potato leaf diseases on imbalanced datasets, with Grad-CAM used to explain the model’s predictions. [87] introduced a novel saliency-based XAI method using perturbation techniques for object detection, which iteratively refines saliency maps to enhance the interpretability of the applied ResNet model while maintaining high accuracy in classifying potato diseases. Finally, [101] presented a tomato health monitoring system that integrates YOLOv8 for detection and MobileNetV3 for real-time counting and classification of diseases, with Grad-CAM++ used to explain the model’s predictions. It is important to mention that CAM-based technology contributed to model verification and highlighted regions with particular texture and color patterns.
A set of research targeted tree and fruit diseases. [75] applied CNN models to detect grape diseases using the PlantVillage dataset, validated with Grad-CAM. [60] introduced GLD-Det, a MobileNet-based model for detecting guava leaf diseases, confirmed by Grad-CAM for real-time mobile applications. This is an important example of a CAM-based solution in real-time prediction, improving the sample prediction explanation. [51] explored the interpretability of CNN models like VGG, GoogLeNet, and ResNet for fruit leaf classification, demonstrating the superior performance of ResNet with Grad-CAM for feature visualization. In another example, [69] integrated a novel module into CNN architectures for fine-grained crop disease classification, with Grad-CAM confirming the model’s focus on relevant features.
@@
Some studies extended the use of XAI techniques beyond agriculture. [62] applied Grad-CAM to validate an ML method for classifying mercury exposure in fish, supporting food safety beyond agriculture. [63] introduced EffiNet-TS, a model based on EfficientNetV2, incorporating an NN to reconstruct images that highlight key symptoms, thereby clarifying the decision-making process. [59] proposed a customized EfficientNetB4 model for high-precision classification of chill leaf diseases, validating the model using Grad-CAM. Similarly, [58] evaluated the performance of four CNN models, with EfficientNetB4 performing best on a dataset of diseased and healthy plant leaves, confirming the models’ focus on critical disease features like rust pustules through Grad-CAM. [73] introduced a meta-learning approach for plant disease detection, interpreted with Task Activation Mapping, a CAM-based technique specifically developed for this study. [61] developed a convolutional ensemble network using lightweight CNNs like MobileNetV2, validated by Grad-CAM. Furthermore, [40] employed a Vision Transformer model for plant disease classification, with Grad-CAM confirming the model’s focus on relevant disease features.
@@
There are numerous applications of AI in the field of fruit quality, particularly those involving the use of XAI to enhance trust in model predictions. An important case is presented by [97], who proposed an ensemble learning framework for fruit plant disease detection using multiple deep learning models, incorporating LIME across all models as an additional tool for result evaluation. [86] proposed a method for classifying various banana diseases—including Cordana, Black Sigatoka, Pestalotiopsis, and Fusarium Wilt—by analyzing leaf images using EfficientNetB0 and employing Grad-CAM to enhance classification accuracy and interpretability. [91] introduced an interpretable AI-based method for localizing mildew symptoms in grapevine using EfficientNetV2S and Grad-CAM. [93] presented LEViT, a Vision Transformer model for tree leaf disease classification, incorporating Grad-CAM to ensure reliable and interpretable results. An example highlighting the need to apply multiple XAI techniques for reliable results is [94], who developed an AI-based system for date palm classification—capable of identifying diseases and assessing fruit ripeness—using VGG16 in combination with SHAP, LIME, Grad-CAM, and Grad-CAM++. [100] proposed a modified MobileNetV2-based model to enhance the classification of cucumber leaf diseases, ensuring result explainability through the integration of LIME. [95] aimed to improve the explainability of deep learning models—specifically a ResNet50 model—used in citrus disease detection, by introducing a novel model-agnostic, local explainer for image-based classification called the Multi-objective Genetic Algorithm Explainer (MOGAE). [96] introduced a CNN-based approach for detecting mulberry leaf diseases, utilizing the MobileNetV3Small model and Grad-CAM to align model predictions with expert assessments.
@@
More recent applications focus on crops, the primary source of human sustenance, highlighting the growing role of AI and XAI in ensuring food security [70] developed MaizeNet, a CNN framework combining clustering for maize crop image segmentation and classification. Grad-CAM was applied to explain the model, providing severity assessments and crop loss estimation. [65] employed CNNs to quantify rice grain chalkiness caused by high nighttime temperatures, using Grad-CAM to localize affected areas. [57] proposed a convolution-based method for rice disease detection, with Grad-CAM highlighting the model’s effectiveness even in complex scenarios. [81] proposed a novel DL model that combines DenseNet for feature extraction with an SVM for classifying healthy and diseased sugarcane plants, incorporating LIME to enhance trust and usability. [82] and [98] applied LRP to enhance VGG16 models for identifying crop leaf diseases, aiming to improve performance. [89] developed a deep transfer learning-based framework for diagnosing rice leaf diseases, leveraging various DL models and integrating Grad-CAM to enhance the system’s reliability for farmers. [90] incorporated LIME into an EfficientNet-based model to address trust issues in plant disease classification. Since Maize Streak Disease poses a serious threat to maize crops, [92] introduced a CNN-based framework for its diagnosis, incorporating SHAP and LIME. Finally, [80] proposed a comparative framework integrating Bayesian optimization for hyperparameter tuning across CNN-based models—InceptionNet, MobileNet, ResNet, and RegNet—to diagnose rice plant diseases, leveraging LIME to enhance the interpretability of model behaviour.
@@
4.2 Mixed Explanation
@@
Regarding tabular data, [24] introduces a novel model for classifying pistachio species by combining feature selection, XAI-based interpretation with LIME, and classification with LR with 90.090.090.090.0%.
@@ 5 Explaining Authenticity and Traceability
By addressing the authenticity and traceability of the food supply chain, we identified a wider application of XAI techniques. This area emerged as the second most significant food-related task application of XAI. Table 2 summarizes the works surveyed in this section.
Table 2: Summary of the works introducing applying XAI for authenticity and traceability surveyed in Section 5, according to their data type and explanation type (labelled as “Expl. type”).
@@
[122] Spectral Visual
[123, 124] Tabular Visual
[33, 125, 126, 127, 28, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145] Tabular Numerical
[23] Time series Numerical
@@
5.1 Visual Explanation
Recent studies using pictorial data have advanced the variety traceability and authenticity verification of agricultural products. [115] developed a CNN model for herb variability identification, using Grad-CAM to highlight relevant herb parts while ignoring background noise. [114] focused on maize seed classification with a ResNet model, while [116] applied HSI and deep learning to classify hybrid okra variability seeds. More recently, [120] proposed the application of various CNN models to classify fungal species, followed by the use of Grad-CAM to interpret the model predictions
@@ 5.1 Visual Explanation
5.2 Numerical Explanation
@@
Several studies have applied advanced ML techniques using tabular data for crop yield prediction, integrating multiple data sources and employing SHAP for interpretability. [127] demonstrated the effectiveness of using SHAP with an AI model for tabular data analysis in aeroponics through data fusion from multiple sensors. Similarly, [125] used XGBoost and SVM to analyze factors affecting rice production, validating model decisions with LIME. [28] applied XGBoost for soybean yield prediction, with SHAP highlighting key factors such as near-infrared light and temperature. [126] further explored soybean yield estimation, emphasizing the role of the vegetation index using SHAP.
@@
Some studies incorporated satellite and meteorological data for improved predictions. [129] utilized LSTM trained on multisource data, applying Integrated Gradients and SHAP to identify critical factors like enhanced vegetation index and temperature. [128] examined the impact of extreme weather on crop yields, revealing sensitivity differences among crops and regions.
@@
Soil water content has also been a focus of ML models in agricultural management. [131] introduced TPE-CatBoost, incorporating soil moisture and environmental factors, with SHAP demonstrating model sensitivity to environmental changes. [130] used TPE-GBDT to map soil water content across the Yellow River Delta, identifying key variables such as soil texture and vegetation. [33] applied SVMs and SHAP to highlight essential factors in digital soil mapping, reinforcing the integration of terrain and geological data for effective agricultural management. The idea of selecting the most suitable soil has also been explored by [142], who aimed to improve crop quality by classifying different soil types using an ML model, and applied SHAP to highlight the most important features influencing the model’s decisions. Similarly, [136] presented an RF model for predicting soil fertility, using SHAP to highlight various physicochemical soil properties that determine fertility levels.
@@
There is also a substantial body of work focused on the traceability and analysis of environmental conditions to enhance the production and quality of crops such as rice, wheat, and maize, leveraging SHAP or LIME to identify the most influential features utilized by the models in performing the given tasks. [132] proposed an ML model for crop prediction, integrating Genetic Algorithms for hyperparameter optimization and RF for classification, while applying XAI techniques such as LIME and SHAP to enhance classifier interpretability—ultimately supporting farmers in optimizing agricultural planning, reducing crop losses, and improving productivity. [133] presented ML models for crop classification and yield prediction, leveraging XAI techniques such as LIME and Feature Importance to enhance model interpretability. Similarly, [141] aimed to provide accurate crop yield predictions by using generative algorithms to optimize a DNN, and employed LIME to explain the model’s outputs. [144] proposed a method for selecting optimal crops based on environmental and soil conditions, utilizing Radial Basis Functions and SHAP. [135] introduced XAI-CROP, an ML-based crop recommendation system improved by including LIME to explain predictions, designed to assist farmers in selecting optimal crops by analyzing soil characteristics, historical crop performance, and weather patterns. A similar tool was developed by [134], who employed various ML models to recommend optimal crops for specific regions, analyzing the results using LIME and SHAP. [139] aimed to enhance the interpretability of AI-driven crop yield predictions by integrating saliency maps and SHAP analysis into KNN models. [138] leveraged an XGBoost model combined with SHAP values to map and understand the influence of weather and soil variables on wheat yield in Eastern Australia. [140] introduced ML-based regression methods along with XAI techniques—SHAP and LIME—to predict crop yields and assess the impact of climate change on agriculture.
@@
Finally, several studies propose applications similar to those previously discussed, but adapted to different food products. In particular, [137] applied ML models—specifically tree-based ensemble methods—and LIME to classify blackcurrant powders based on image texture features. [143] proposed using ML-based models, such as Random Forest and SHAP, to enhance coffee quality assessment—traditionally reliant on subjective evaluation—by contributing to the standardization of coffee grading. [145] examined the integration of ANN and XAI techniques, such as Feature Importance, to enhance quality control strategies in the agri-food industry, with a specific focus on milk quality classification.
5.3 Rule-based Explanation
Authenticity and traceability have not been deeply explored in rule-based explanations. [23] highlighted the need to monitor low-cost, automated, and interpretable irrigation systems using time series data. To address this, they proposed a new system called Vital, which integrates IoT sensors, a data management platform, and a fuzzy rule-based decision support system to automate irrigation. The system was evaluated through pilot cases and effectively automated the irrigation process, monitoring and managing open-field installations that provided water.
5.4 Mixed Explanation

### v2

@@ 4 Explaining Food Safety
Table 3: Summary of the works introducing applying XAI for Food Safety topic surveyed in Section 4, according to their data type and explanation type (labelled as “Expl. type”).
@@
[198, 57] Tabular Mixed
4.1 Visual Explanation for the Food Safety topic
Pictorial data.
Numerous studies using pictorial data have focused on detecting plant diseases in staple crops such as maize, rice, and wheat, underscoring the importance of accurate disease identification in the food supply. [41] developed a transfer learning methodology enhancing MobileNetV2 with CAM to diagnose plant diseases in maize and rice. [44] and [81] used CNN models to detect maize and peanut diseases, applying channel attention and pruning techniques to improve feature extraction. Similarly, [64] applied CNNs with Grad-CAM to distinguish healthy from infected wheat, effectively identifying disease-affected areas, while [90] used similar techniques to extend crop disease classification to a resource-constrained setting. [150] introduced C-DenseNet, a modified CNN model, to grade wheat stripe rust severity, validated using Grad-CAM++ [39]. [71] developed a YOLOv5s-based model with MobileNetV3 and C3Ghost modules to detect fusarium head blight (FHB) in wheat, using Grad-CAM. In addition, to address the complexities of varying disease images, [42] improved MobileNetV2 with a Location-wise Soft Attention mechanism and CAM, demonstrating its practical utility for identifying crop diseases under diverse conditions.
In contrast, some studies explored alternative crops and agricultural sectors. [166] used high-resolution video data and a CNN-based object detection model to monitor pecan tree health, focusing on xylella disease, validated by Grad-CAM to highlight critical canopy features. [274] introduced T-RNet, a Transformer-Embedded ResNet model, for cassava leaf disease detection, demonstrating focus on relevant areas through Grad-CAM visualizations. [6] proposed LeafVisionNet, a new CNN model integrated with SHAP for black gram leaf disease detection, designed to operate in resource-constrained agricultural settings. A similar need to operate in environments with limited data was addressed by [223], who proposed a cost-effective, explainable framework with DenseNet and various XAI techniques, including SmoothGrad, Grad-CAM++, and Score-CAM, for betel leaf disease detection. [129], [29], and [30] aimed to develop a non-destructive and explainable image-based method for identifying adulteration in red chili powder. [129] proposed using a GAN to increase the dataset size and a feature selection model to improve the training of AI models, with SHAP and LIME employed to validate the model’s learning. [29] and [30] used two DenseNet-based models, leveraging Grad-CAM and LIME to explain the results.
@@ 4.1 Visual Explanation for the Food Safety topic
Similarly, [75] and [148] applied LIME and Grad-CAM to a DenseNet-based model developed for classifying tomato leaf diseases. Moreover, [96] proposed the use of a MobileNetV2 model combined with data augmentation and reweighting techniques for accurate classification of potato leaf diseases on imbalanced datasets, with Grad-CAM used to explain the model’s predictions. [22] introduced a novel saliency-based XAI method using perturbation techniques for object detection, which iteratively refines saliency maps to enhance the interpretability of the applied ResNet model while maintaining high accuracy in classifying potato diseases. Two other significant works have been proposed by [14] and [53], who introduce two new lightweight CNN architectures, XSE-TomatoNet and XLTLDisNet, designed to detect diseases in tomato plants. In both studies, XAI plays a crucial role, utilizing various techniques such as SHAP, LIME, Grad-CAM, and Grad-CAM++. The comparison of these explanations helps validate the developed models. Finally, [185] presented a tomato health monitoring system that integrates YOLOv8 for detection and MobileNetV3 for real-time counting and classification of diseases, with Grad-CAM++ used to explain the model’s predictions. It is important to mention that CAM-based technology contributed to model verification and highlighted regions with particular texture and color patterns.
A set of research targeted tree and fruit diseases. [142] applied CNN models to detect grape diseases using the PlantVillage dataset, validated with Grad-CAM. [157] introduced GLD-Det, a MobileNet-based model for detecting guava leaf diseases, confirmed by Grad-CAM for real-time mobile applications. This is an important example of a CAM-based solution in real-time prediction, improving the sample prediction explanation. [249] explored the interpretability of CNN models like VGG, GoogLeNet, and ResNet for fruit leaf classification, demonstrating the superior performance of ResNet with Grad-CAM for feature visualization. In another example, [270] integrated a novel module into CNN architectures for fine-grained crop disease classification, with Grad-CAM confirming the model’s focus on relevant features. [239] proposed DynLeafNet, a dynamic, computationally efficient, and explainable CNN-based architecture for automatic plant disease recognition, trained on the PlantVillage dataset, with results explained using Grad-CAM++.
@@
Some studies extended the use of XAI techniques beyond agriculture. [144] applied Grad-CAM to validate an ML method for classifying mercury exposure in fish, supporting food safety beyond agriculture. [72] introduced EffiNet-TS, a model based on EfficientNetV2, incorporating an NN to reconstruct images that highlight key symptoms, thereby clarifying the decision-making process. [183] proposed a customized EfficientNetB4 model for high-precision classification of chill leaf diseases, validating the model using Grad-CAM. Similarly, [203] evaluated the performance of four CNN models, with EfficientNetB4 performing best on a dataset of diseased and healthy plant leaves, confirming the models’ focus on critical disease features like rust pustules through Grad-CAM. [43] introduced a meta-learning approach for plant disease detection, interpreted with Task Activation Mapping, a CAM-based technique specifically developed for this study. [40] developed a convolutional ensemble network using lightweight CNNs like MobileNetV2, validated by Grad-CAM. Similarly, [159] implemented MobileNetV2, integrating Grad-CAM and LIME for explainability, to develop an end-to-end explainable system for maize disease diagnosis that can be deployed in the field. [7] applied SHAP and LIME to explain the results of a CNN developed for early maize leaf disease detection, suitable for mobile deployment. Furthermore, [143] employed a Vision Transformer model for plant disease classification, with Grad-CAM confirming the model’s focus on relevant disease features.
@@
There are numerous applications of AI in the field of fruit quality, particularly those involving the use of XAI to enhance reliability in model predictions. An important case is presented by [169], who proposed an ensemble learning framework for fruit plant disease detection using multiple DL models, incorporating LIME across all models as an additional tool for result evaluation. [13] proposed a method for classifying various banana diseases—including Cordana, Black Sigatoka, Pestalotiopsis, and Fusarium Wilt—by analyzing leaf images using EfficientNetB0 and employing Grad-CAM to enhance classification accuracy and interpretability. [84] introduced an interpretable AI-based method for localizing mildew symptoms in grapevine using EfficientNetV2S and Grad-CAM. [182] presented LEViT, a Vision Transformer model for tree leaf disease classification, incorporating Grad-CAM to ensure reliable and interpretable results. An example highlighting the need to apply multiple XAI techniques for reliable results is [106], who developed an AI-based system for date palm classification—capable of identifying diseases and assessing fruit ripeness—using VGG16 in combination with SHAP, LIME, Grad-CAM, and Grad-CAM++. [215] proposed a modified MobileNetV2-based model to improve the classification of cucumber leaf diseases, ensuring result explainability through the integration of LIME. Cucumbers were also analyzed by [189], who proposed using a CNN to examine plant leaf images for disease diagnosis, leveraging Grad-CAM, Grad-CAM++, and Eigen-CAM to explain the results. [161] aimed to improve the explainability of DL models—specifically a ResNet50 model—used in citrus disease detection, by introducing a novel model-agnostic, local explainer for image-based classification called the Multi-objective Genetic Algorithm Explainer (MOGAE). [197] introduced a CNN-based approach for detecting mulberry leaf diseases, utilizing the MobileNetV3Small model and Grad-CAM to align model predictions with expert assessments. Finally, to enhance crop management by detecting potential diseases, [225] used a combination of two versions of InceptionNet and ResNet to classify diseases in orange plants, explaining the results with LIME, Grad-CAM, and Score-CAM, and comparing the obtained outcomes.
@@
More recent applications focus on crops, the primary source of human sustenance, highlighting the growing role of AI and XAI in ensuring food security [113] developed MaizeNet, a CNN framework combining clustering for maize crop image segmentation and classification. Grad-CAM was applied to explain the model, providing severity assessments and crop loss estimation. [244] employed CNNs to quantify rice grain chalkiness caused by high nighttime temperatures, using Grad-CAM to localize affected areas. [258] proposed a convolution-based method for rice disease detection, with Grad-CAM highlighting the model’s effectiveness even in complex scenarios. [66] proposed a novel DL model that combines DenseNet for feature extraction with an SVM for classifying healthy and diseased sugarcane plants, incorporating LIME to enhance reliability and usability. [226] and [174] applied LRP to enhance VGG16 models for identifying crop leaf diseases, aiming to improve performance. [186] developed a deep transfer learning-based framework for diagnosing rice leaf diseases, leveraging various DL models and integrating Grad-CAM to enhance the system’s reliability for farmers. [163] incorporated LIME into an EfficientNet-based model to address reliability issues in plant disease classification. Since Maize Streak Disease poses a serious threat to maize crops, [102] introduced a CNN-based framework for its diagnosis, incorporating SHAP and LIME. Finally, [238] proposed a comparative framework integrating Bayesian optimization for hyperparameter tuning across CNN-based models—InceptionNet, MobileNet, ResNet, and RegNet—to diagnose rice plant diseases, leveraging LIME to enhance the interpretability of model behaviour.
Spectral data.
@@
4.2 Numerical Explanation for the Food Safety topic
Tabular data.
Recent studies have introduced numerical explanations to increase the transparency of AI applications for addressing food safety issues. In particular, all the analyzed works used SHAP to identify the most significant features contributing to the models’ outputs. All of these studies relied on tabular data. [19] proposed a scalable and interpretable framework that combined biomonitoring, spatial screening, and CatBoost to assess health risks from airborne organic pollutants affecting honey, demonstrating that honey, together with XAI-enabled ML, can provide actionable and transparent evidence for environmental health management. [123] developed a scalable and interpretable ML framework based on RF that simultaneously predicted daily milk yield and identified health risks in large dairy herds by integrating environmental, nutritional, and physiological data. [8] proposed a hybrid XAI-based feature selection framework that combines Permutation Feature Importance (PFI) and SHAP values to identify the most informative handcrafted features extracted from insect images for classification. PFI and SHAP are used to quantify and rank feature importance, and the final reduced feature set is obtained by taking the intersection of the top-ranked features from both methods, leading to improved classification accuracy. A work closely related to [8] was presented in [173], where the authors extracted handcrafted texture and intensity features from fish images to develop an accurate ML-based system for early fish disease diagnosis.
4.3 Mixed Explanation for the Food Safety topic
Pictorial data.
@@ 4.3 Mixed Explanation for the Food Safety topic
Tabular data.
Regarding tabular data, [198] introduced a novel model for classifying pistachio species by combining feature selection, XAI-based interpretation with LIME, and classification with LR. Finally, [57] developed a Decision Support System using an ensemble of ML models to predict the risk of wheat powdery mildew in winter wheat crops before visible symptoms appear, thereby supporting proactive disease management in agriculture. The results were subsequently explained using various XAI techniques, including SHAP and LIME.
@@ 5 Explaining Authenticity and Traceability
By addressing the Authenticity and Traceability challenges of the food supply chain, we identified a wider application of XAI techniques. This area emerged as the second most significant food-related task application of XAI. Table 4 summarizes the works surveyed in this section.
Table 4: Summary of the works introducing applying XAI for the Authenticity and Traceability topic surveyed in Section 5, according to their data type and explanation type (labelled as “Expl. type”).
@@
[252, 255] Tabular Visual
[97] Spectral Visual
[277, 3, 120, 233, 88, 83, 141, 267, 263, 133, 17, 112, 206, 63, 38, 184, 69, 111, 155, 135, 24, 201, 158, 1, 65, 181, 209] Tabular Numerical
[119, 194, 269] Time series Numerical
[18, 220] Spectral Numerical
@@
5.1 Visual Explanation for the Authenticity and Traceability topic
Pictorial data.
Recent studies using pictorial data have advanced the variety traceability and authenticity verification of agricultural products. [273] developed a CNN model for herb variability identification, using Grad-CAM to highlight relevant herb parts while ignoring background noise. [256] focused on maize seed classification with a ResNet model, while [264] applied HSI and DL to classify hybrid okra variability seeds. More recently, [188] proposed the application of various CNN models to classify fungal species, followed by the use of Grad-CAM to interpret the model predictions
@@ 5.1 Visual Explanation for the Authenticity and Traceability topic
5.2 Numerical Explanation for the Authenticity and Traceability topic
Tabular data.
@@ 5.2 Numerical Explanation for the Authenticity and Traceability topic
Several studies have applied advanced ML techniques using tabular data for crop yield prediction, integrating multiple data sources and employing SHAP for interpretability. [233] demonstrated the effectiveness of using SHAP with an AI model for tabular data analysis in aeroponics through data fusion from multiple sensors. [181] proposed a CNN-based framework for predicting maize yield, focusing on augmenting the often-limited datasets and explaining the results through SHAP and LIME. Similarly, [3] used XGBoost and SVM to analyze factors affecting rice production, explaining model decisions with LIME. [88] applied XGBoost for soybean yield prediction, with SHAP highlighting key factors such as near-infrared light and temperature. [120] further explored soybean yield estimation, emphasizing the role of the vegetation index using SHAP.
@@
Some studies incorporated satellite and meteorological data for improved predictions. [141] utilized Long Short-Term Memory (LSTM), a type of DNN, trained on multisource data, applying Integrated Gradients and SHAP to identify critical factors like enhanced vegetation index and temperature. [83] examined the impact of extreme weather on crop yields, revealing sensitivity differences among crops and regions.
@@
Soil water content has also been a focus of ML models in agricultural management. [263] introduced TPE-CatBoost, incorporating soil moisture and environmental factors, with SHAP demonstrating model sensitivity to environmental changes. [267] used TPE-GBDT to map soil water content across the Yellow River Delta, identifying key variables such as soil texture and vegetation. [277] applied SVMs and SHAP to highlight essential factors in digital soil mapping, reinforcing the integration of terrain and geological data for effective agricultural management. The idea of selecting the most suitable soil has also been explored by [24], who aimed to improve crop quality by classifying different soil types using an ML model, and applied SHAP to highlight the most important features influencing the model’s decisions. Similarly, [38] presented an RF model for predicting soil fertility, using SHAP to highlight various physicochemical soil properties that determine fertility levels. [209] proposed an IoT framework that collected soil and weather data to provide real-time crop recommendations, using RF and a NN as predictive models and explaining the results with LIME. Finally, [65] used XGBoost to analyze soil data and predict wheat varieties in Morocco, once again explaining the results with SHAP and PDP.
@@
There is also a substantial body of work focused on the traceability and analysis of environmental conditions to enhance the production and quality of crops such as rice, wheat, and maize, leveraging SHAP or LIME to identify the most influential features utilized by the models in performing the given tasks. [133] proposed an ML model for crop prediction, integrating Genetic Algorithms for hyperparameter optimization and RF for classification, while applying XAI techniques such as LIME and SHAP to enhance classifier interpretability—ultimately supporting farmers in optimizing agricultural planning, reducing crop losses, and improving productivity. [17] presented ML models for crop classification and yield prediction, leveraging XAI techniques such as LIME and Feature Importance to enhance model interpretability. Similarly, [135] aimed to provide accurate crop yield predictions by using generative algorithms to optimize a DNN, and employed LIME to explain the model’s outputs. [158] proposed a method for selecting optimal crops based on environmental and soil conditions, utilizing Radial Basis Functions and SHAP. [206] introduced XAI-CROP, an ML-based crop recommendation system improved by including LIME to explain predictions, designed to assist farmers in selecting optimal crops by analyzing soil characteristics, historical crop performance, and weather patterns. A similar tool was developed by [112], who employed various ML models to recommend optimal crops for specific regions, analyzing the results using LIME and SHAP. [111] aimed to enhance the interpretability of AI-driven crop yield predictions by integrating saliency maps and SHAP analysis into KNN models. [69] leveraged an XGBoost model combined with SHAP values to map and understand the influence of weather and soil variables on wheat yield in Eastern Australia. [1] also introduced a tool for predicting subsurface nutrient levels, specifically nitrogen, potassium, and phosphorus, in cabbage cultivation by using plant growth characteristics to train an NN and explaining the obtained results with SHAP and LIME. [155] introduced ML-based regression methods along with XAI techniques—SHAP and LIME—to predict crop yields and assess the impact of climate change on agriculture.
@@
Finally, several studies propose applications similar to those previously discussed, but adapted to different food products. In particular, [184] applied ML models—specifically tree-based ensemble methods—and LIME to classify blackcurrant powders based on image texture features. [201] proposed using ML-based models, such as RF and SHAP, to enhance coffee quality assessment—traditionally reliant on subjective evaluation—by contributing to the standardization of coffee grading. [63] examined the integration of ANN and XAI techniques, such as Feature Importance, to enhance quality control strategies in the agri-food industry, with a specific focus on milk quality classification.
Time series data.
Some studies use time series as the data type for training AI models. In these cases, all the works employ explanations based on feature importance, calculated by permuting the model’s learned features and assessing how these variations affect the model’s output. [119] developed a method based on a Graph CNN model to predict grain temperatures across different ecological zones, grain types, and granary structures, supporting early warning and management. [269] embedded mathematical equations into a DL model to control the training process and the features learned from soil data, with the aim of predicting soil moisture.
Spectral data.
Two studies instead used spectral data to address authenticity and traceability tasks, both employing SHAP to provide numerical explanations of their models. [18], using NIR spectra, built a rapid, non-destructive, and interpretable pipeline to evaluate Astragali Radix quality by origin and antioxidant activity. The pipeline consisted of a classifier to authenticate the root’s origin, followed by an SVM regression model to determine its antioxidant activity. Similarly, [220] aimed to determine the origin and grade of green tea from elemental fingerprints, using Inductively Coupled Plasma Atomic Emission Spectrometry to collect data from different tea varieties and classifying them with a DNN.
5.3 Mixed Explanation for the Authenticity and Traceability topic
Pictorial data.